Solve using simplex method

z = 3000x + 5000y

2x + y <= 16

x + 2y <= 11

x + 3y <= 15


1
Expert's answer
2021-11-18T09:45:08-0500

Max "z=3000 x+5000 y" subject to "2 x+y \leq 16 x+2 y \leq 11\ and\ x, y \geq 0"

The problem is converted to canonical form by adding slack, surplus and artificial variables as appropriate

1. As the constraint- 1 is of type ' "\leq" ' we should add slack variable "S_{1}"

2. As the constraint- 2 is of type ' "\leq" ' we should add slack variable "S_{2}"

After introducing slack variables

"\operatorname{Max} z=3000 x+5000 y+0 S_{1}+0 S_{2}"

subject to

 "\begin{array}{rr}

2 x+y+S_{1} & =16 \\

x+2 y+S_{2} & =11

\end{array}"

 

and "x, y, S_{1}, S_{2} \geq 0"



Since all "Zj-Cj\u22650"

Hence, optimal solution is arrived with value of variables as :

"x=7,y=2"

Max "z=31000"